OpenStudy (anonymous):

(y+x)(ax+b) yax + yb + ax^2 + xb y^2 + a^2 + x^4 + b^2 Is this right?

OpenStudy (anonymous):

Please help me, I did the work I just need someone to check it

OpenStudy (trojanpoem):

Multiply each term times each term in the second parentheses. (y+x)(ax+b) = y * ax + y * b + x * ax + x * b = yax + yb + ax^2 + xb As easy as that.

OpenStudy (anonymous):

Great, thats what I got. Thank you